# Reflected generalized BDSDEs driven by non-homogeneous Lévy processes and obstacle problems for stochastic integro-PDEs with nonlinear Neumann boundary conditions

### v2

@@ 4.4 The obstacle problem of SIPDE-NBC
In this final segment, we report our primary finding, extending the previous results to formulate a probabilistic formula for an obstacle problem of SIPDE-NBC.
Initially, we shall incorporate the subsequent assumption on the barrier.
  • $(\mathcal{H}.6)$
@@
Using the same arguments as in [37], we get the following continuity result:
@@
Utilizing the results from Section 4.3 and employing the methodology outlined in [1] (Theorem 4.2, pp.1168), we derive:
Theorem 31.
Under conditions $(\mathcal{H}.1)$$(\mathcal{H}.6)$, the random field $u(\omega,t,x)=Y^{t,x}_{t}$ is stochastic viscosity solution for the subsequent obstacle problem of SIPDE-NBC
$\left\{\begin{array}[]{lll}\displaystyle{\big(u(t,x)-\mathcal{S}(t,x)\big)\wedge\left(-\frac{\partial u}{\partial t}(t,x)-\mathcal{L}_{t}u(t,x)-f\left(t,x,u(t,x),\left(u^{1}_{k}(t,x)\right)^{\infty}_{k=1}\right)\right)}\\ \hskip 256.0748pt\displaystyle+g\left(t,x,u(t,x)\right)dB_{t}=0,&(t,x)\in[0,T[\times G\\ \\ \displaystyle u(T,x)=H(x),&x\in\bar{G}\\ \displaystyle\frac{\partial u}{\partial n}(t,x)+h(t,x,u(t,x))=0,&x\in\partial G,\end{array}\right$.
@@ Appendix A Special generalized backward doubly SDEs with Lipschitz drivers
The existence and uniqueness results for generalized BDSDEs with a non-homogeneous Lévy process (GBDSDE-NL) are provided in this section for the specific case where the coefficients $f$ and $g$ are Lipschitz and depend only on $y$, while the driver $g$ does not depend on the variables $(y,z)$. We first consider the case of standard generalized BDSDE-NL without reflection. Then, using these results, we extend the analysis to the reflected case with a continuous obstacle $S$.

### v3

@@ 4.4 The obstacle problem of SIPDE-NBC
First, let us consider the following assumption on the barrier.
  • $(\mathcal{H}.5)$
@@
Using the same arguments as in [36], we get the following continuity result.
@@
With the help of the result obtained in Section 4.3 and by the same method used in [1] (Theorem 4.2, pp.1168), we establish the existence of a stochastic viscosity solution to the following obstacle problem of SIPDE-NBC:
$\left\{\begin{array}[]{lll}\displaystyle{\big(u(t,x)-\mathcal{S}(t,x)\big)\wedge\left(-\frac{\partial u}{\partial t}(t,x)-\mathcal{L}_{t}u(t,x)-f\left(t,x,u(t,x),\left(u^{1}_{k}(t,x)\right)^{\infty}_{k=1}\right)\right)}\\ \\ \hskip 199.16928pt\displaystyle+g\left(t,x,u(t,x)\right)dB_{t}=0,&(t,x)\in[0,T[\times G\\ \\ \displaystyle u(T,x)=H(x),&x\in\bar{G}\\ \\ \displaystyle\frac{\partial u}{\partial n}(t,x)+h(t,x,u(t,x))=0,&x\in\partial G,\end{array}\right$. (4.69)
@@ Appendix A Special generalized backward doubly SDEs with Lipschitz drivers
The existence and uniqueness results for generalized BDSDEs with jumps and a non-homogeneous Lévy process (GBDSDE-NL) are provided in this section for the specific case where the coefficients $f$ and $g$ are Lipschitz and depend only on $y$, while the driver $g$ does not depend on the variables $(y,z)$. We first consider the case of standard generalized BDSDE-NL without reflection. Then, using these results, we extend the analysis to the reflected case with a continuous obstacle $S$.